ADNAN MUHAMMED

Grade 12,

The orbital angular momentum corresponding to n=4 and m=-3 is :

The orbital angular momentum corresponding to n=4 and m=-3 is :

Grade:12

1 Answers

Sudhir Bishnoi
6 years ago
Orbital angular momentum= √[l(l+1)] ×(h÷2π)
Given n=4 therefore l can be 0,1,2,3
But m=-l to +l
Therefore  for l=0, m=0
for l=1, m=0,-1,1
for l=2, m=-2,-1,0,1,2
for l=3, m=-3,-2,-1,0,1,2,3
Therefore for formula l=3
Putting l=3 in formula we get oam=3.64×10-34 J.s